OpenStudy (anonymous):

314=-n + 576 Show work please

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

sorry another go 314 - 576 = -n -262 = -n n = 262

OpenStudy (anonymous):

haha k thanks

OpenStudy (anonymous):

try these and practise rearranging equations as much as you can.

OpenStudy (anonymous):

How do u do -1/4w - 1=6

OpenStudy (anonymous):

-1/4w - 1 = 6 -1/4w = 6+1 -1/4w =7 w = 7/(-1/4) w = -28 should be close to being correct, post up on main page to let someone double check it.

OpenStudy (anonymous):

thanks